OpenStudy (kainui):

I need help with this algebra problem, you will definitely have to use the product log.

OpenStudy (kainui):

\[\Large a^p-a=np\] I need to solve for p

OpenStudy (kainui):

True, but I thought it would be fun to do it without using WA just as an exercise. I want to figure out the steps with anyone who's interested.

OpenStudy (kainui):

Why am I interested in this? Well for all prime numbers there is an n and a you can plug in to get p.

OpenStudy (kainui):

Haha yes, I just thought it would be interesting to solve for the prime number in it. =P

OpenStudy (anonymous):

this is fermat theorem , i thought i posted this comment :O

OpenStudy (anonymous):

it says a^p=a mod p for any p prime s.t GCD(a,p)=1

OpenStudy (kainui):

I guess alternatively, how do I get wolfram alpha to graph this: http://www.wolframalpha.com/input/?i=f%28x%2Cy%29%3D%28-y+W%28-x%5E%28-x%2Fy%29ln%28x%29%2Fy%29-xlnx%29%2F%28ylnx%29

ganeshie8 (ganeshie8):

gcd(a,p) = 1 is not needed

OpenStudy (anonymous):

w is wronsckin ?

ganeshie8 (ganeshie8):

but looks Kainui wants the continuous version

ganeshie8 (ganeshie8):

W is product log function, you want to abuse wronskians als here ha!

ganeshie8 (ganeshie8):

@dan815 @UnkleRhaukus might know plotting this in matlab

OpenStudy (kainui):

Since this is true for any integer a, we can just set it equal to 2. And I believe this will map all the prime numbers uniquely 1-to-1 to another integer. \[\Large \frac{2^p-2}{p}=n\] This seems pretty cool.

OpenStudy (anonymous):

ganesh gcd is needed else it would be a^p=0 mod p

OpenStudy (anonymous):

that work kai , but note n,a are changeable

OpenStudy (anonymous):

u can change 2 by any other prime xD but its means nothing

OpenStudy (anonymous):

eh , im not expert with w function though i wud like to learn it

OpenStudy (kainui):

Here are two sites that really helped me, there are some good exercises on there! It's really fun, I'll help you if you have questions. =D http://www.had2know.com/academics/lambert-w-function-calculator.html https://luckytoilet.wordpress.com/tag/product-log/

OpenStudy (kainui):

True, we can make it anything, but I figured I'd make it 2 since it's the smallest possible one to make computation easier.

OpenStudy (watchmath):

\(a,n\) are real numbers?

OpenStudy (kainui):

No, they are natural numbers.

OpenStudy (kainui):

To anyone curious looking for patterns, I've listed the primes on the left and to the right are different values of n for consecutive integers starting at 2. \[\Large \frac{a^p-a}{p}=n\] So if you look in the third row we have p=5, and it lists for a=2, n=6; for a=3, n=48; for a=4, n=204; etc... ``` 2: 1, 3, 6, 10, 15, 21, 28, 36, 3: 2, 8, 20, 40, 70, 112, 168, 240, 5: 6, 48, 204, 624, 1554, 3360, 6552, 11808, 7: 18, 312, 2340, 11160, 39990, 117648, 299592, 683280, 11: 186, 16104, 381300, 4438920, 32981550, 179756976, 780903144, 2852823600, 13: 630, 122640, 5162220, 93900240, 1004668770, 7453000800, 42288908760, 195528140640, 17: 7710, 7596480, 1010580540, 44878791360, 995685849690, 13684147881600, 132458812569720, 981010688215680, 19: 27594, 61171656, 14467258260, 1003867701480, 32071565263710, 599941851861744, 7585009898729256, 71097458824894312, 23: 364722, 4093181688, 3059510616420, 518301258916440, 34336096654504472, 1189945536525257216, 29: 18512790, 2366564736720, 9938978487990060, 6422914307692954624, 31: 69273666, 19924948267224, 148764065110560896, 37: 3714566310, 12169835294351280, 41: 53634713550, 889585277491970432, 43: 204560302842, 7633882962663652352, 47: 2994414645858, 53: 169947155749830, 57: 2528336632909752, ```

OpenStudy (watchmath):

So what are you looking for? finding \(p\) given \(a\) and \(n\) (which is not always exist) or finding \(n\) given \(a\) and \(p\) which is what you did on the table?

OpenStudy (kainui):

What I'm saying is we have this function: \[\Large f(n) = \frac{-W(\frac{1}{\sqrt[n]{4}}\ln \frac{1}{\sqrt[n]{2}})}{\ln2}-\frac{2}{n}\] and on the domain where n is only natural numbers, the only values in the range of f(n) that are integers are all prime numbers.

OpenStudy (kainui):

To put another way to make it apparent how cool this is! \[domain \ f = \mathbb{N} \]\[(range \ f) \cap \mathbb{N} = set\ of\ all\ primes\]

OpenStudy (kainui):

Basically the way I know it is just like it says on this site: https://www.math.hmc.edu/funfacts/ffiles/10007.5.shtml

OpenStudy (anonymous):

i see , thanks for the new thing to think about xD